OpenStudy (narissa):

Can someone help me solve for x--1/7=3/4x-1

OpenStudy (anonymous):

1/7= 3/(4x-1) cross multiply (4x-1) 1=7*3 4x-1=21 4x=22 x=22/4 =5.5

OpenStudy (narissa):

but the answer has to be a fraction

OpenStudy (thatonegirl_):

11/2

OpenStudy (thatonegirl_):

\[5 \frac{ 1 }{2 }\]

OpenStudy (thatonegirl_):

multiply the denominator by the whole number, then add that by the numerator. Put it all over 2. 5*2=10. 10+1=11. 11/2

OpenStudy (thatonegirl_):

put it all over the denominator* woops